OpenStudy (anonymous):

A pitcher throws a baseball straight into the air with a velocity of 72 feet/sec. If acceleration due to gravity is −32 ft/sec2, how many seconds after it leaves the pitcher's hand will it take the ball to reach its highest point? Assume the position at time t = 0 is 0 feet.

OpenStudy (perl):

You can use the equation: $$ \Large { y = -\frac g 2 t^2 + v_ot + y_o }$$

OpenStudy (perl):

v_0 = velocity at time zero y_o = height at time zero g = acceleration due to gravity

OpenStudy (anonymous):

i dont know the last two vairables

OpenStudy (anonymous):

@perl

OpenStudy (perl):

we are given that v_o = 72 ft/s, and y_0 = 0

OpenStudy (perl):

You can use the equation: $$ \Large { y = -\frac g 2 t^2 + v_ot + y_o \\ \therefore \\ y = -\frac{32}{2}\cdot t^2 + 72t + 0 }$$

OpenStudy (anonymous):

y=-16t^2+72t

OpenStudy (anonymous):

@perl

OpenStudy (anonymous):

@rational

OpenStudy (perl):

yes thats correct. now find the maximum of that . that occurs when t = -b / (2a)

OpenStudy (anonymous):

i dont know what those values are. @perl